The average price for personal computers has recently decreased from $300 to $250. As a result,...

  1. The average price for personal computers has recently decreased from $300 to $250. As a result, the quantity of personal computers bought has increased by 20 percent. What would be the price elasticity of demand for personal computers?
Answer #1

Answer : % changes in price = [(250 - 300) / 300] * 100

=> % changes in price = - 0.17 * 100 = - 17%

% changes in quantity demanded = 20%

Price elasticity of demand (Ed) = % changes in quantity demanded / % changes in price level

=> Ed = 20 / (-17)

=> Ed = - 1.2

Therefore, here the price elasticity of demand is - 1.2 .